# Learning Cross-Task Generalities Across Graphs via Task-trees

### v1

@@ 3 Task-Trees Serve As Basic Learning Instances on Graphs
The proof can be found in Appendix D.3. This theorem outlines key factors affecting model generalization on downstream tasks, such as the transferability of task-trees (${\mathcal{C}}_{\delta}({\mathcal{E}}_{\mathcal{P}}(g,\phi))^{\delta}$) and the quality of the pretrained encoder (${\mathcal{E}}_{\mathcal{P}}(g,\phi)$). With regard to the number of task-trees, we find that while increasing the number of fine-tuning samples contributes to more stable optimization ($\frac{4{\mathcal{C}}_{1}}{n}\sqrt{\sum_{i=1}^{n}\|\phi(T_{i})\|^{2}}$), it does not significantly reduce the generalization bound ($2\sqrt{\frac{\log(1/v)}{n}}$). This provides theoretical evidence that a reasonable number of fine-tuning samples can be sufficient for training a model with strong generalization capabilities. Moreover, the discrepancy between the pretraining and fine-tuning distributions ($\sum_{x\in{\mathcal{X}}_{\phi}}\|{\mathcal{T}}_{\phi}(x)-{\mathcal{P}}_{\phi}(x)\|$) is crucial—smaller distribution gaps lead to better generalization. This highlights the importance of increasing the diversity of pretraining data, which provides a boarder pretraining distribution ${\mathcal{P}}$. It also supports the potential of developing specialized models for specific domains based on a pretrained general model, which is discussed in Section 4.2.
4 GIT: Graph Generality Identifier on Task-Trees
@@
The theoretical analysis demonstrates the feasibility of building graph foundation models based on task-trees. In this section, we apply these theorems to develop a cross-task, cross-domain GFM called GIT, with the aim of empirically validating the Task-Tree Generality Assumption.
4.1 General Model: Pretraining to Acquire General Knowledge
We propose a task-tree reconstruction task as a pretext for pretraining. The key idea is to use two corrupted task-trees to reconstruct each other, thereby capturing corruption-invariant semantics of the task-trees. Given a set of task-trees $\{T^{t}_{1},...,T^{t}_{n}\}$ sampled from a graph database [^1] We assume the node attributes of all graphs have been aligned using methods such as LLMs (Liu et al., 2024) or SVD (Zhao et al., 2024a). While proper node attribute alignment is important for GFMs, it is beyond the scope of this paper. We use text-attributed graphs in our experiments due to the data availability (Chen et al., 2024b; Zhang et al., 2024b; Feng et al., 2024), and use Sentence Bert (Reimers & Gurevych, 2019) to convert node features of all graphs into 768-dim vectors, following Liu et al. (2024)., we apply corruption techniques to generate two views of each task-tree, denoted as $\{\hat{T}^{t}_{1},...,\hat{T}^{t}_{n}\}$ and $\{\tilde{T}^{t}_{1},...,\tilde{T}^{t}_{n}\}$. For corruption, we use random edge masking and random attribute masking, as proposed by Zhu et al. (2020), due to its computational efficiency. We then use an encoder $\phi$ to obtain embeddings for the corrupted task-trees, resulting in $\{\hat{{\bm{z}}}_{1},...,\hat{{\bm{z}}}_{n}\}$ and $\{\tilde{{\bm{z}}}_{1},...,\tilde{{\bm{z}}}_{n}\}$. Note that the task-tree embedding is defined as the average of the embeddings of task-relevant nodes, avoiding the introduction of inductive biases that could lead the model to learn incorrect patterns. The loss function is
${\mathcal{L}}=\underbrace{\frac{1}{2n}\sum_{i=1}^{n}\Bigl{[}\bigl{\|}\rho(g(\hat{{\bm{z}}}_{i}))-\operatorname{sg}[\rho(\tilde{{\bm{z}}}_{i})]\bigr{\|}^{2}+\bigl{\|}\rho(g(\tilde{{\bm{z}}}_{i}))-\operatorname{sg}[\rho(\hat{{\bm{z}}}_{i})]\bigr{\|}^{2}\Bigr{]}}_{\text{Reconstruction}}+\underbrace{\sum_{i=1}^{n}D_{\mathrm{KL}}({\bm{h}}\|{\bm{z}}_{i})}_{\text{Reg.}},$ (6)
@@ 4.1 General Model: Pretraining to Acquire General Knowledge
where $g$ is a non-linear MLP projector, $\rho({\bm{z}})=({\bm{z}}/\|{\bm{z}}\|)$ serves for normalization, $\operatorname{sg}$ is the stop-gradient operation, and ${\bm{h}}$ is the average of all instances ${\bm{z}}$. The reconstruction loss captures the semantics of the task-trees in a predictive manner, while the regularizer ensures the embeddings are projected into a shared space by minimizing the KL divergence between individual instances and their center. Additional analysis is provided in Appendix C.
4.2 Specialized Model: Specification via Instruction Tuning
@@
where $\psi$ is the instruction generation function for each task-tree, and $\kappa$ is the corresponding loss function. In this paper, as we use text-attributed graphs in our experiments, we define instructions as the embeddings of label descriptions encoded by a LLM, which is similar to Liu et al. (2024), and we use mean squared error as the loss function $\kappa$.
@@ 5 Experiment
Figure 3: The model performance on all datasets in the in-context setting.
Table 1: We report the model performance across five graph domains: academia, e-commerce, knowledge base, molecular, and temporal graphs, with results averaged over all graphs within each domain. Note that -G and -S represent the general and specialized versions of GIT, respectively. The comprehensive results with additional baselines and ablation studies can be found in Appendix F.
@@ 5.1 Experimental Setup
The detailed experimental settings are provided in Appendix E, and the comprehensive experimental results and analysis are presented in Appendix F. In the following, we briefly introduce the datasets, baselines, and basic evaluation protocols.
Datasets. Our experiments are based on text-attributed graphs due to data availability. Specifically, we include over 30 graphs spanning five domains: academic networks, e-commerce networks, knowledge graphs, molecular graphs, and temporal graphs. Detailed information is provided in Appendix E.1. For model pretraining, we use the citation network Arxiv, the e-commerce network Products, knowledge graphs WN18RR and FB15K237, and molecular graphs Chemblpre and PCBA. For supervised fine-tuning (SFT) during specialization, we use Arxiv for academic networks, Products for e-commerce networks, FB15K237 for knowledge graphs, and PCBA for molecular networks. For temporal graphs, which are e-commerce temporal graphs, we also use Products for SFT to assess robustness under temporal distribution shifts.
Baselines. We employ a wide range of baselines, including supervised GNNs (Kipf & Welling, 2017; Veličković et al., 2018; Xu et al., 2019), self-supervised GNNs such as BGRL (Thakoor et al., 2022) and GraphMAE (Hou et al., 2022), graph foundation models like OFA (Liu et al., 2024), and domain-specific expert models. These include Ultra (Galkin et al., 2024) for knowledge graphs and KVPLM (Zeng et al., 2022), MoMu (Su et al., 2022), Galactica (Taylor et al., 2022), and GIMLET (Zhao et al., 2023) for molecular graphs. Further details are provided in Appendix E.2.
Experimental Protocols. We use GraphSAGE (Hamilton et al., 2017) as the encoder, and each experiment is repeated at least five times with different seeds to reduce the impact of randomness. We evaluate three settings: fine-tuning, in-context learning, and zero-shot learning. Fine-tuning is the standard approach where the model is fine-tuned on downstream tasks. In-context learning is a few-shot learning method without fine-tuning, where we randomly select $k$ samples from a given class, average these samples to form prototypes, and use the prototypes for classification. Following Liu et al. (2024); He & Hooi (2024), we randomly sample 500 5-way 3-shot tasks. If the number of classes is fewer than 5, the number of ways is set to the number of classes. Zero-shot learning operates similarly, but instead of using sample averages, it employs class description embeddings encoded by an LLM as the prototypes for prediction. Considering the metrics, we use accuracy in node classification and edge classification, and AUC in graph classification and link prediction.
@@ 5.2 Main Results
Results on all graphs. We present the model performance across all graphs for both the basic fine-tuning and in-context learning settings in Figure 4 and Figure 3, respectively. Results for zero-shot learning are provided in Appendix F.1. In these figures, we compare the performance of GIT [^2] We report the best performance across general GIT and specialized GIT. and the best baselines. Note that we do not include domain expert models here since they are only applicable to specific domains. We observe that GIT consistently outperforms the best baselines across most datasets, highlighting the advantages of treating task-trees as the fundamental learning instance and providing empirical validation of the Task-Tree Generality Assumption.
Figure 4: The model performance on all datasets in the fine-tuning setting.
Table 2: In-context learning results on knowledge graphs, where Ultra is the domain expert. We bold the best results and underline the second-best. The results with std. are in Table 20 and 21.
Domain Expert Ultra 60.69 82.45 74.35 75.98 90.22 33.89 41.37 89.29 68.53
General Model GraphMAE 55.20 61.41 54.30 61.01 86.42 32.43 31.58 70.15 56.56
OFA 55.27 55.14 50.20 62.40 88.41 30.23 34.94 79.15 56.97
GIT - S 57.90 77.19 72.14 76.99 90.80 34.85 42.02 90.49 67.80
Table 3: Zero-shot results on molecule graphs with domain experts KVPLM (Zeng et al., 2022), MoMu (Su et al., 2022), Galactica (Taylor et al., 2022), and GIMLET (Zhao et al., 2023).
HIV BBBP BACE TOXCAST CYP450 TOX21 MUV PCBA Avg.
Domain Expert KVPLM 61.20 60.20 51.26 50.96 59.22 49.17 61.72 48.11 55.23
MoMu 50.26 49.81 66.56 52.38 57.98 57.57 60.51 51.50 55.82
Galactica-1.3B 33.85 53.94 56.48 51.23 46.86 49.46 57.15 52.02 50.12
GIMLET 66.24 59.39 69.57 59.04 71.25 61.19 64.39 62.11 64.15
General Model BGRL 55.27 53.72 33.74 49.00 60.99 46.40 39.90 42.39 47.68
GraphMAE 46.48 49.08 30.76 48.22 60.55 49.17 48.17 45.10 47.19
OFA 47.96 50.61 34.35 49.70 61.96 52.73 52.48 54.14 50.49
GIT - S 66.14 62.16 52.27 58.30 69.75 63.45 65.32 65.26 62.83
@@
Results by domains. For a clearer understanding, we present the experimental results averaged across each domain in Table 1, with additional results and extra baselines provided in Appendix F, categorized by domain. In this table, we compare model performance across three settings: zero-shot learning, 3-shot in-context learning, and basic fine-tuning, alongside Supervised GNN (Sup. GNN), self-supervised GraphMAE (Hou et al., 2022), and the graph foundation model OFA (Liu et al., 2024). The “held-out avg.” indicates the average performance on all graphs, excluding those used in pretraining and specialization, to eliminate their influence on downstream tasks. We observe that even the general GIT significantly outperforms the baselines in some settings. After specialization, the specialized GIT further boosts performance, particularly in zero-shot and in-context learning settings, which supports our theoretical findings that post-training a pretrained model on domain-specific task-trees can make the model more suitable for specific domains.
The analysis of specialization. More encouragingly, applying the proposed specialization method to other graph learning models also improves their performance in specific domains (Table 17). This highlights the potential of post-training to enhance the capabilities of GNNs. To further investigate the specialization process, we conducted experiments to assess the impact of SFT data (Figure 12 and Table 24). The results show that SFT data can improve the performance of the general model across most settings. However, different SFT datasets have varying effects on different downstream tasks, underscoring the need for better SFT data selection methods. Additionally, we observe that the specialized model may suffer from reduced general inference capability, similar to the specialization tax seen in LLMs (Zhang et al., 2023). It is crucial to develop techniques that balance GIT’s specialized and general capabilities.
5.3 Case Studies by Domains
Molecule Graphs. We present the zero-shot graph classification results in Table 3. Following Zhao et al. (2023), we evaluate zero-shot performance on the original testing set. In addition to comparing with general algorithms applicable to all graphs, like GraphMAE, and OFA, we include domain-expert models designed specifically for molecules, such as KVPLM (Zeng et al., 2022), MoMu (Su et al., 2022), Galactica (Taylor et al., 2022), and the recent SOTA GIMLET (Zhao et al., 2023). Notably, our specialized GIT model achieves the second-best average performance following GIMLET, and even surpasses GIMLET on 4 out of 8 datasets, while significantly outperforming other general models and domain experts. This observation supports our theoretical analysis, demonstrating that post-training a pretrained model with domain-specific instances can effectively build a specialized model for that domain.
Knowledge Bases. We present the in-context edge classification results on knowledge graphs in Table 2, comparing our GIT model with general models and the domain expert Ultra (Galkin et al., 2024). The specialization is done using FB15K237. We observe that our specialized GIT significantly outperforms the baselines, even approaching the performance of Ultra, a foundation model specifically designed for knowledge graphs. Even more impressively, GIT surpasses Ultra on 5 out of 8 datasets, highlighting the potential of using task-trees as basic learning instances and aligning with our theoretical analysis.
Figure 5: Comparison between GIT and best baselines on academia networks.
Table 4: The link prediction results on academia networks.
Cora Citeseer Pubmed Arxiv23 DBLP Arxiv Avg.
Best Baseline 87.34 ± 0.88 87.52 ± 0.98 89.83 ± 0.35 91.45 ± 0.44 98.29 ± 0.07 97.50 ± 0.08 90.96
Academic Networks. We present the node classification and link prediction results on academic networks in Figure 5 and Table 4, respectively. We compare the performance of GIT specialized via Arxiv, general GIT, and the best baseline. Both versions of GIT consistently outperform the best baselines across datasets and settings, with the specialized version generally performing better than the general one. For link prediction, we randomly split the train/val/test sets in a 70%/15%/15% ratio. GIT also achieves the best performance in this setting. Moreover, specialization consistently improves model performance, even though the process is optimized at the node level rather than the edge level, demonstrating the potential for cross-task transferability through specialization.
5.4 Comparison Between Task-Trees and Subgraphs
Instances Subgraph Task-tree
@@ 5.4 Comparison Between Task-Trees and Subgraphs
Table 5: Performance comparison between methods with different learning instances.
Unlike our proposed task-trees, some approaches use $k$-hop subgraphs extracted from graphs as the basic learning units (Huang et al., 2023; Liu et al., 2024). In node classification, for example, ego-graphs are extracted around each node, and the label of the central node is assigned to the induced subgraph, transforming node classification into subgraph classification. A similar process can be applied to convert edge-level and graph-level tasks into subgraph-level tasks. This approach involves (1) extracting ego-graphs around task-relevant nodes and (2) applying GNNs to learn graph-level embeddings for classification. However, this extraction process introduces additional computational overhead, increasing both time consumption and memory usage due to the need to store and process the induced subgraphs. Furthermore, the information within these subgraphs is not always easily learned by basic GNNs, as they may struggle to capture essential substructures needed for learning graph-level embeddings (Garg et al., 2020; Chen et al., 2020; Zhang et al., 2024a), thereby limiting the effectiveness of using subgraphs as basic learning instances.

### v2

@@ 3 Why Do Task-Trees Serve as Basic Learning Instances?
The proof can be found in Appendix D.3. This theorem outlines key factors affecting model generalization on downstream tasks, such as the transferability of task-trees (${\mathcal{C}}_{\delta}({\mathcal{E}}_{\mathcal{P}}(g,\phi))^{\delta}$) and the quality of the pretrained encoder (${\mathcal{E}}_{\mathcal{P}}(g,\phi)$). With regard to the number of task-trees, we find that while increasing the number of fine-tuning samples contributes to more stable optimization ($\frac{4{\mathcal{C}}_{1}}{n}\sqrt{\sum_{i=1}^{n}\|\phi(T_{i})\|^{2}}$), it does not significantly reduce the generalization bound ($2\sqrt{\frac{\log(1/v)}{n}}$). This provides theoretical evidence that a reasonable number of fine-tuning samples can be sufficient for training a model with strong generalization capabilities. Moreover, the discrepancy between the pretraining and fine-tuning distributions ($\sum_{x\in{\mathcal{X}}_{\phi}}\|{\mathcal{T}}_{\phi}(x)-{\mathcal{P}}_{\phi}(x)\|$) is crucial—smaller distribution gaps lead to better generalization. This highlights the importance of increasing the diversity of pretraining data, which provides a boarder pretraining distribution ${\mathcal{P}}$. It also supports the potential of developing specialized models for specific domains based on a pretrained general model, discussed in Section 4.2.
4 Graph Generality Identifier on Task-Trees
@@
The theoretical analysis demonstrates the feasibility of building graph foundation models based on task-trees. We apply these theorems to develop our GIT model, with the aim of empirically validating the Task-Tree Generality Assumption.
As our focus is to align task spaces on graphs, we adopt existing text-attributed graph benchmarks (Chen et al., 2024b; Zhang et al., 2024b; Feng et al., 2024) to enable feature space alignment. In particular, we use Sentence Bert (Reimers & Gurevych, 2019) to convert node features of all graphs into 768-dim vectors, following Liu et al. (2024).
4.1 GIT-G: Pretraining to Acquire General Knowledge
We propose a task-tree reconstruction task as a pretext for pretraining. The key is to use two corrupted task-trees to reconstruct each other, thereby capturing corruption-invariant semantics. Given a set of task-trees $\{T^{t}_{1},...,T^{t}_{n}\}$ sampled from a graph database, we apply corruption techniques to generate two views of each task-tree, denoted as $\{\hat{T}^{t}_{1},...,\hat{T}^{t}_{n}\}$ and $\{\tilde{T}^{t}_{1},...,\tilde{T}^{t}_{n}\}$. For corruption, we use random edge masking and random attribute masking (Zhu et al., 2020) due to its computational efficiency. We then use an encoder $\phi$ to obtain embeddings for the corrupted task-trees, resulting in $\{\hat{{\bm{z}}}_{1},...,\hat{{\bm{z}}}_{n}\}$ and $\{\tilde{{\bm{z}}}_{1},...,\tilde{{\bm{z}}}_{n}\}$. Inspired by (Thakoor et al., 2022), we perform reconstruction as
${\mathcal{L}}$ $=\frac{1}{2n}\sum_{i=1}^{n}\Bigl{[}\|\rho(g(\hat{{\bm{z}}}_{i}))-\operatorname{sg}[\rho(\tilde{{\bm{z}}}_{i})]\|^{2}$
$+\|\rho(g(\tilde{{\bm{z}}}_{i}))-\operatorname{sg}[\rho(\hat{{\bm{z}}}_{i})]\|^{2}\Bigr{]}+\sum_{i=1}^{n}D_{\mathrm{KL}}({\bm{h}}\|{\bm{z}}_{i}),$ (7)
@@ 4.1 GIT-G: Pretraining to Acquire General Knowledge
where $g$ is a non-linear MLP projector, $\rho({\bm{z}})=({\bm{z}}/\|{\bm{z}}\|)$ serves for normalization, $\operatorname{sg}$ is the stop-gradient operation, and ${\bm{h}}$ is the average of all instances ${\bm{z}}$. The reconstruction loss captures the semantics of the task-trees in a predictive manner, while the KL regularizer ensures the embeddings are projected into a shared space by minimizing the KL divergence between individual instances and their center.
4.2 GIT-S: Specification via Instruction Tuning
@@
where $\psi$ is the instruction generation function for each task-tree, and $\kappa$ is the corresponding loss function. In this paper, as we use text-attributed graphs in our experiments, we define instructions as the embeddings of label descriptions encoded by a LLM, which is similar to Liu et al. (2024), and we use mean squared error as the loss function $\kappa$. Additional model analysis is provided in Appendix B.
@@ 5 Experiment
Table 2: We report the model performance across five graph domains: academia, e-commerce, knowledge base, molecular, and temporal graphs, with results averaged over all graphs within each domain. Note that -G and -S represent the general and specialized versions of GIT, respectively. The comprehensive results can be found in Appendix F.
@@
Figure 3: The model performance on all datasets in the fine-tuning setting. For GIT, the best result is selected between GIT-G and GIT-S, while the best baseline performance is chosen among GCN/GAT/GIN, BGRL, GraphMAE, and OFA.
@@ 5.1 Experimental Setup
Datasets. Our experiments are based on text-attributed graphs. Specifically, we include over 30 graphs spanning five domains: academic networks, e-commerce networks, knowledge graphs, molecular graphs, and temporal graphs. Detailed information is provided in Appendix E.1. During pretraining, we use citation network Arxiv, e-commerce network Products, knowledge graphs WN18RR and FB15K237, and molecular graphs Chemblpre and PCBA. During specialization, we use Arxiv for academic networks, Products for e-commerce networks, FB15K237 for knowledge graphs, and PCBA for molecular networks. For temporal graphs, which are e-commerce temporal graphs, we also use Products for SFT to assess robustness under temporal distribution shifts.
Baselines. We employ a wide range of baselines, including supervised GNNs like GCN (Kipf & Welling, 2017), GAT (Veličković et al., 2018), GIN (Xu et al., 2019), self-supervised GNNs such as BGRL (Thakoor et al., 2022) and GraphMAE (Hou et al., 2022), graph foundation models like OFA (Liu et al., 2024). We also include some domain expert models. Details are provided in Appendix E.2.
Experimental Protocols. We use GraphSAGE (Hamilton et al., 2017) as the encoder, and each experiment is repeated at least five times with different seeds. We evaluate three settings: fine-tuning, in-context learning, and zero-shot learning. Fine-tuning is the standard approach where the model is fine-tuned on downstream tasks. In-context learning is a few-shot learning method without fine-tuning, where we randomly select $k$ samples from a given class, average these samples to form prototypes, and use the prototypes for classification. Following Liu et al. (2024); He & Hooi (2024), we randomly sample 500 5-way 3-shot tasks. If the number of classes is fewer than 5, the number of ways is set to the number of classes. Zero-shot learning operates similarly, but instead of using sample averages, it employs class description embeddings encoded by an LLM as the prototypes for prediction. Considering the metrics, we use accuracy in node classification and edge classification, and AUC in graph classification and link prediction. Additional details are provided in Appendix E.
@@ 5.2 Main Results
Domain-wise Results. Table 2 presents the experimental results averaged across each domain, with detailed results provided in Appendix F. The “held-out avg.” represents the average performance on all graphs, excluding those used in pretraining and specialization, ensuring an unbiased evaluation of downstream tasks. Notably, even the general GIT-G model surpasses baseline methods in certain settings. After specialization, the specialized GIT-S achieves additional performance gains, particularly in zero-shot and in-context learning. These findings align with our theoretical analysis, demonstrating that post-training a pretrained model enhances its adaptability to specialized domains.
Dataset-wise Results. Figure 3 illustrates the model performance across all graphs in the standard fine-tuning setting. We compare the performance of GIT against the strongest baseline methods. Notably, GIT consistently outperforms the best baselines across most datasets, further reinforcing the effectiveness of task-trees as learning instances.
Impact of Specialization. We summarize three key observations: (1) Specialization (GIT-S) enhances the performance of GIT-G across most settings (Table 2). However, the impact of specialization varies depending on the dataset used (Appendix Figure 9 and Table 23). (2) Specialization does not significantly degrade model performance on other domains (Table 25). (3) Applying the proposed specialization approach to other models also yields performance improvements in specific domains (Table 3), underscoring the potential of post-training in enhancing the capabilities of GNNs.
Table 3: Ablation study on training strategies, where the performance is averaged across all academic networks.
Base Model Expert Model
0-shot 3-shot Finetune 0-shot 3-shot Finetune
GraphMAE 15.30 51.51 75.57 17.89 55.88 75.26
OFA 14.19 50.15 75.12 17.26 54.88 75.97
GIT 15.36 53.31 75.53 18.38 55.10 75.47
General Model Specialized Model
0-shot 3-shot Finetune 0-shot 3-shot Finetune
GraphMAE 15.42 49.25 73.81 20.31 51.21 74.05
OFA 13.98 45.93 72.18 20.05 46.87 73.04
GIT 14.88 54.00 75.82 23.45 55.18 75.88
Table 4: Performance comparison between methods with different basic learning instances.
Instances Subgraph Tree
@@
5.3 Ablation on Training Strategies
We conduct an ablation study on academic networks to evaluate the impact of different training strategies. Specifically, we examine four approaches: (1) Base Model: Pretraining on the target graph. (2) Expert Model: Pretraining on all academic networks. (3) General Model: Pretraining on the default pretraining datasets. (4) Specialized Model: Pretraining on the default datasets followed by specialization on Arxiv. The results, averaged across all academic graphs, are reported in Table 3. Notably, the general model of GIT maintains stable performance relative to both the base and expert models. In contrast, GraphMAE and OFA exhibit performance degradation when transitioning from the base and expert models to the general model. This finding underscores the potential of task-trees in mitigating such negative transfer. Furthermore, specialization in GIT allows its performance to closely approximate that of expert models trained exclusively on academic graphs.
5.4 Comparison to Domain Experts
We compare GIT to domain experts in molecular and knowledge graphs, as these domains require domain-specific knowledge. Our findings indicate that the specialized GIT (GIT-S) closely approximates the performance of these experts. In molecular graphs, GIT-S achieves an average performance of 62.83, ranking second to the state-of-the-art GIMLET (64.15) (Zhao et al., 2023) while outperforming other domain experts (Zeng et al., 2022; Su et al., 2022; Taylor et al., 2022), whose best performance is 55.82 (Table 22). Similarly, in knowledge graphs, GIT-S attains an average score of 67.80, approaching the KG expert Ultra (68.53) (Galkin et al., 2024), as shown in Table 21 in Appendix.
5.5 Comparison Between Task-Trees and Subgraphs